OpenStudy (anonymous):

show the algebraic inequality. show work. Tammy has the following scores on her business law test 92,99,84,91. what score must she earn in order to an A average. and A at her univ starts at 90

OpenStudy (anonymous):

are you supposed to find out how much she must earn in a 5th test to average 90 or more?

OpenStudy (anonymous):

(92+99+84+91+x)/5 >= 90 92+99+84+91+x >= 450 x >= 450 - 366 x >= 84